# CREStE: Scalable Mapless Navigation with Internet Scale Priors and Counterfactual Guidance

### v1

@@ IV-A2 Reward Function Model Architecture
We implement $r_{\phi}$ using a 0.5M parameter Multi-Scale Fully Convolutional Network (MS FCN) first used by Wulfmeier et al. [50]. We select this model architecture as it enforces spatial invariance and considers features at multiple scales.
@@ IV-B CREStE Training Procedure
To train CREStE, we first optimize $\Theta$ and freeze the parameters before training $r_{\phi}$ using our active reward learning framework with counterfactuals. Altogether, our full learning objective is:
@@
where $\mathcal{L}_{\text{rgbd}}$ and $\mathcal{L}_{\text{bev}}$ supervise $\Theta$ and $\mathcal{L}_{\text{IRL}}$ supervise $r_{\phi}$. Sec. IV-B1 breaks down the objective terms $\mathcal{L}_{\text{rgbd}}$ and $\mathcal{L}_{\text{bev}}$ and how we leverage VFMs to generate training labels. Sec. IV-B2 defines our counterfactual-based IRL objective $\mathcal{L}_{\text{IRL}}$, procedure for obtaining counterfactual annotations, and theoretical derivation for $\mathcal{L}_{\text{IRL}}$.
@@ IV-B1 Training the Perceptual Encoder $\Theta$
To train $\Theta$, we propose a distillation label generation module that leverages large VFMs to generate training labels. Our labels enable us to learn representations that are robust to perceptual aliasing, spatially grounded, and rich in navigation priors. As shown in Fig. 2, we generate semantic feature maps $\hat{y}_{\text{semantic}}$ and completed depth labels $\hat{y}_{\text{depth}}$ to distill semantic and geometric priors to $f_{\text{rgbd}}$. Then, we generate BEV map labels for static and dynamic entities ($\hat{y}_{\text{bev, static}}$, $\hat{y}_{\text{bev, dynamic}}$) and BEV elevation map labels $\hat{y}_{\text{bev, elev}}$ to distill entity priors and enforce robustness to lifting artifacts when training $f_{\text{bev}}$. Next, we present our distillation label generator and training procedure for $f_{\text{rgbd}}$ and $f_{\text{bev}}$.
Distillation Label Generator. Fig. 2 visually depicts the inputs and outputs for our label generator. Using sequential SE(3) robot poses and synchronized RGB–point cloud pairs ($o_{\text{rgb, 1:t}}$, $o_{\text{cloud, 1:t}}$), our Distillation Label Generator produces five training labels for supervising $\Theta$. Below, we detail the label generation procedure for a single timestep t, separating it into two parts: (i) labels supervising $f_{\text{rgbd}}$, and (ii) labels supervising $f_{\text{bev}}$.
i) Generating Training Labels for the RGB-D Encoder $f_{\text{rgbd}}$. We generate $\hat{y}_{\text{semantic}}$ by passing $o_{\text{rgb, t}}$ through a frozen Dinov2 encoder and bilinearly interpolating the spatial dimension of our output feature map to match the spatial resolution of $y_{\text{semantic}}$. We generate $\hat{y}_{\text{depth}}$ by projecting $o_{\text{cloud, t}}$ to the image and applying bilateral filtering [31] for edge-aware inpainting, producing a dense depth map that respects object boundaries.
ii) Generating Training Labels for the BEV Inpainting Backbone $f_{\text{bev}}$. To generate $\hat{y}_{\text{bev, dynamic}}$, we prompt SAM2 with bounding box detections from a set of commonly encountered dynamic categories (e.g. vehicles, pedestrians) to obtain dynamic entity labels. We follow the approach in Osep. et al. [29], backprojecting dynamic labels to 3D using the corresponding point cloud $o_{\text{cloud, t}}$ and clustering the points at multiple density thresholds using DBSCAN [7]. We retain the DBSCAN clusters that exceed a minimum intersection-over-union (IoU) overlap with the entity-labeled clusters and project the matched points to the dynamic entity BEV map.
To generate $\hat{y}_{\text{bev, static}}$, we first obtain static entity labels by prompting SAM2 with a grid of query points (generating dynamic entity labels via bounding box queries as before), mask out overlapping regions between grid-queried and dynamic labels to isolate the static masks in each frame, and then apply an iterative greedy merging strategy that fuses overlapping masks (above an IoU threshold) across consecutive frames — treating unmatched IDs as new entities. Finally, we project and accumulate these entity-consistent multi-frame labels in BEV using known robot poses to produce the final static entity BEV map. To see our algorithmic formulation describing this procedure, please refer to the Appendix.
To generate $\hat{y}_{\text{bev, elev}}$, we accumulate static entity-labeled 3D points across sequential frames using robot poses (generating static entity labels as before), assign each 3D point to a grid cell, and compute each cell’s minimum elevation by averaging the N lowest 3D points that fall into that cell.
Training the RGB-D Encoder $f_{\text{rgbd}}$. We jointly train $f_{\text{rgbd}}$ via backpropagation from $f_{\text{semantic}}$ and $f_{\text{depth}}$. First, the semantic decoder head $f_{\text{semantic}}$ is supervised via an L2/MSE loss that minimizes the error between $y_{\text{semantic}}$ and $\hat{y}_{\text{semantic}}$. Second, the depth completion head $f_{\text{depth}}$ is supervised using a cross-entropy classification loss $\mathcal{L}_{\mathcal{C}\mathcal{E}}$, where $\hat{y}_{\text{depth}}$ is uniformly discretized into bins. Empirically, this captures depth discontinuities better than regression [34]. Altogether the training objective for the RGB-D backbone is:
@@
Training the BEV Inpainting Backbone $f_{\text{bev}}$. We train $f_{\text{bev}}$ with three losses, one for each BEV decoder head, that leverage $\hat{y}_{\text{bev, static}}$, $\hat{y}_{\text{bev, dynamic}}$, and $\hat{y}_{\text{bev, elev}}$ to distill entity priors, mitigate projection artifacts from noisy depth predictions, and teach our model to reason about occluded areas. We train $f_{\text{bev, static}}$ and $f_{\text{bev, dynamic}}$ with Supervised Contrastive Loss [18] ($\mathcal{L}_{\text{contrastive}}$) using BEV map labels $\hat{y}_{\text{bev, static}}$ and $\hat{y}_{\text{bev, dynamic}}$. $\mathcal{L}_{\text{contrastive}}$ optimizes the embedding space such that features of the same entity remain close while repelling those from different entities. This is key to learning continuous feature maps from discrete BEV map labels. We train $f_{\text{bev, elev}}$ to predict $\hat{y}_{\text{bev, elev}}$ using $l_{1}$ regression loss. Our full training objective for $f_{\text{bev}}$ is:
@@
where $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are tunable hyperparameters. Trained together, $\mathcal{L}_{\text{rgbd}}$ and $\mathcal{L}_{\text{bev}}$ guide the network to learn robust semantic, geometric, and entity-level representations, even for regions that are partially or fully hidden from direct view. For specific hyperparameter settings, please refer to Appendix Sec. X-A.
@@ IV-B2 Training the Reward Function $r_{\phi}$
Even with perceptual representations that encode relevant features for navigation, prior works [8, 15] demonstrate that is difficult to identify the most salient features and how they influence operator preferences. This problem is exacerbated for real-world deployments, where limited expert demonstrations further reduce generalizability. To improve policy learning under limited expert demonstrations, we propose an active learning framework that supplements expert demonstrations with counterfactual annotations obtained via offline operator feedback. Our approach teaches policies to learn rewards that simultaneously penalize suboptimal behavior while rewarding expert-aligned behavior, thereby improving sample efficiency, generalizability, and interpretability.
More formally, suboptimal trajectories enable learning more discriminative reward functions that map scene features to their scalar utilities. Let us define a state-action visitation distribution to be the discounted probability of reaching a state $s$ under a policy $\pi$ and taking action $a$: $\rho^{\pi}(s,a)=(1-\gamma)\sum\gamma^{t}p(s_{t}=s,a_{t}=a|\pi)$. We define each state $s$ to be an xy location on the local BEV grid and our actions $a$ along the 8 connected grid. Overall, our objective for reward learning, referred to as Counterfactual IRL, is simple: For each BEV observation $y_{bev}$, given state-actions sampled from the expert’s visitation distribution $\rho^{E}$ and state-actions sampled from the suboptimal visitation distribution $\rho^{S}$; we obtain a reward function and policy by solving the following optimization problem:
@@
where $\rho_{\pi}$ denotes current policy visitation and $\alpha$ is a tunable hyperparameter that balances the relative importance between suboptimal and expert demonstrations. Using the relationship defined in Eq. 10, we can rewrite Eq. 6 in terms of the state-action visitation distribution to obtain our Counterfactual IRL training objective:
@@
Intuitively, the above objective learns a reward function such that the difference between the expert’s return and agent policy’s return is minimized while ensuring suboptimal counterfactuals have a low return. Assuming non-trivial rewards, which can be enforced using gradient regularization techniques [22], our objective effectively leverages expert and counterfactual demonstrations to learn more discriminative rewards. Next, we describe our active learning framework for learning $r_{\phi}$ from counterfactuals and how we generate these counterfactual annotations. We conclude by deriving the reward learning objective using the Bradley-Terry model of preferences and show connections to Inverse Reinforcement Learning (IRL).
Active Reward Learning from Counterfactuals. We introduce an active reward learning framework that leverages counterfactual demonstrations to teach our reward function $r_{\phi}$ to map features from our BEV feature map $y_{\text{bev}}$ to operator-aligned scalar rewards. Fig. 3 illustrates this framework, which consists of three phases that we shall describe next:
@@
Phase I: Warmstart We first set $\alpha$ equal to zero in $\mathcal{L}_{\text{IRL}}$, effectively training $r_{\phi}$ using only expert demonstrations. This provides a base policy to use for the next phase.
Phase II: Synthetic Counterfactual Generation For each training sample, we infer rewards using $r_{\phi}$ from Phase I and plan trajectories using the same start and goal as the expert trajectory. For samples where the learned and expert policies diverge, we generate alternate trajectories from the start to the goal and prompt the operator to identify which alternate trajectories are counterfactuals (i.e exhibit unwanted behavior).
@@
Phase III: Counterfactual Reward Alignment We retrain $r_{\phi}$ using $\mathcal{L}_{\text{IRL}}$ with expert and counterfactual annotations. We set $\alpha$ to be nonzero to balance the relative importance of expert and counterfactual demonstrations. We repeat phases II and III using the current $r_{\phi}$ until $r_{\phi}$ aligns with operator preferences.
@@
Generating Counterfactual Annotations. To obtain counterfactual annotations for a training sample t in our dataset, we first require the expert trajectory $\Gamma^{E}_{t}$, $o_{\text{rgb, t}}$, and $o_{\text{depth, t}}$ for the sample. We uniformly sample a handful of ”control” states along $\Gamma^{E}_{t}$, excluding the start and goal states. Then, we randomly perturb these control states before planning a kinematically feasible path from the start and goal states such that it reaches all of the perturbed control states. Practically, we implement this using Hybrid A* [6] - however, any kinematic planner will suffice. We generate a handful of alternate trajectories in this manner and prompt the operator to identify the suboptimal trajectories given $o_{\text{rgb, t}}$ and $o_{\text{depth, t}}$. We use the selected suboptimal trajectories to retrain $r_{\phi}$ in our active reward learning framework. For additional implementation details regarding generating alternate trajectories, we refer readers to Appendix Sec. X-B.
@@
Counterfactual IRL Derivation $\mathcal{L}_{\text{IRL}}$. In this section, we derive $\mathcal{L_{\text{IRL}}}$, our IRL objective that leverages expert and counterfactual demonstrations. For context, IRL methods [53, 26] allow for learning reward functions $r_{\phi}$, parameterized by $\phi$, given expert demonstrations. However, they provide no mechanism to incorporate suboptimal trajectories. Suboptimal trajectories are easy to obtain and enable a data flywheel for navigation; the BEV observations obtained from expert runs can simply be relabelled with suboptimal of unsafe trajectories offline without any more environmental interactions. Motivated by this idea, we derive $\mathcal{L_{\text{IRL}}}$, a general and principled way to learn from suboptimal and expert trajectories jointly under a single objective.
@@
This property extends to any suboptimal visitations, and as a consequence of linearity of expectations, to any convex combination of the current policy’s visitation and any other suboptimal visitation distribution. Mathematically,
@@
$\min_{\pi}\max_{\phi}\left(J^{E}(r_{\phi})-(\alpha J^{S}(r_{\phi})+(1-\alpha)J^{\pi}(r_{\phi})\right).$ (14)
@@ V-1 Global and Local Planning and Controls
Our global and local planning modules work together to identify the next local subgoal for local path planning. Given a user-specified GPS end goal $G_{N}$, we use OpenStreetMap [27] to obtain a semi-dense sequence of coarse GPS goals $G=[G_{1},...,G_{N}]$ spaced 10 meters apart. To select the next GPS subgoal, we compute the set of distances $D=\{||g-G_{N}||\text{ }|\text{ }g\in G\cup\{G_{t}\}\}$, where $D$ contains the distance of the robot $G_{t}$ from $G_{N}$ and the distance of each GPS subgoal in $G$ from $G_{N}$. From the set of subgoals with a smaller distance to $G_{N}$ than $G_{t}$, we select the farthest subgoal to use as the next subgoal. We project this subgoal on the edge of the local planning horizon, a 6 meter circle around the robot, giving us a carrot for local path planning.
We adopt a DWA [9] style approach to local path planning, where we enumerate a set of constant curvature arcs (31 in our case) from the egocentric robot frame. We compute learned cost for each trajectory by sampling points along each each arc and computing the discounted cost at each point using our predicted reward map. We simply invert and normalize our reward map to the range $[0,1]$ to convert it to a costmap. Finally, we compute the distance between the last point on the arc with the local carrot to obtain a goal-reaching cost. We multiply the learned and goal-reaching costs by tunable weights before selecting the trajectory with the lowest cost. Finally, we perform 1D time optimal control [30] to generate low-level actions to follow this trajectory. Empirically, we find that sampling 30 points and using a discount factor of 0.95 is sufficiently dense. We find that that tuning the goal-reaching cost to be 1/10th the importance of the learned cost achieves good balance between goal-reaching and adhering to operator preferences.
@@
  • ($Q_{3}$) How much do counterfactual demonstrations improve learned policies in challenging urban scenes?

### v2

@@ IV-A2 Reward Function Model Architecture
To ensure our reward function enforces spatial invariance and considers multi-scale features, we implement $r_{\phi}$ using a 0.5M parameter Multi-Scale Fully Convolutional Network (MS FCN) first used by Wulfmeier et al. [47]. We supervise $r_{\phi}$ using our counterfactual IRL loss, which we describe along with our training procedure for $\Theta$ in the next section.
@@ IV-B CREStE Training Procedure
We train CREStE by first optimizing $\Theta$ and freezing the parameters before training $r_{\phi}$ using our counterfactual-based active reward learning framework. Altogether, our full learning objective is:
@@
where $\mathcal{L}_{\text{rgbd}}$ and $\mathcal{L}_{\text{bev}}$ supervise $\Theta$ and $\mathcal{L}_{\text{IRL}}$ supervise $r_{\phi}$. Sec. IV-B1 breaks down the our training and label generation procedure for supervising $\Theta$ and Sec. IV-B2 defines our counterfactual IRL objective $\mathcal{L}_{\text{IRL}}$ and active reward learning framework for training $r_{\phi}$.
@@ IV-B1 Training the Perceptual Encoder $\Theta$
We jointly train $f_{\text{rgbd}}$ via backpropagation from $f_{\text{semantic}}$ and $f_{\text{depth}}$. The semantic decoder head $f_{\text{semantic}}$ is supervised via an MSE loss that minimizes the error between the predicted and ground truth feature maps $y_{\text{semantic}}$ and $\hat{y}_{\text{semantic}}$. The depth completion head $f_{\text{depth}}$ is supervised using a cross-entropy classification loss $\mathcal{L}_{\mathcal{C}\mathcal{E}}$, where the ground truth depth $\hat{y}_{\text{depth}}$ is uniformly discretized into bins. Empirically, this captures depth discontinuities better than regression [31]. Altogether, the training objective for the RGB-D backbone is:
@@
We supervise $f_{\text{bev}}$ with three losses, one for each BEV decoder head, using ground truth static instance maps $\hat{y}_{\text{bev, static}}$, dynamic instance maps $\hat{y}_{\text{bev, dynamic}}$, and elevation maps $\hat{y}_{\text{bev, elev}}$. We use Supervised Contrastive Loss [16] ($\mathcal{L}_{\text{contrastive}}$) for training $f_{\text{bev, static}}$ and $f_{\text{bev, dynamic}}$ to predict continuous feature maps from discrete instance labels $\hat{y}_{\text{bev, static}}$ and $\hat{y}_{\text{bev, dynamic}}$. We train $f_{\text{bev, elev}}$ to predict $\hat{y}_{\text{bev, elev}}$ using $l_{1}$ regression loss. Our full training objective for $f_{\text{bev}}$ is:
@@
where $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are tunable hyperparameters. We refer readers to Appendix Sec. X-B for additional training details.
To scalably obtain training labels for $\Theta$, we design a distillation label generation module that leverages VFMs to automatically generate training labels. As shown in Fig. 2, our label generator only requires sequential SE(3) robot poses and synchronized RGB–point cloud pairs ($o_{\text{rgb, 1:t}}$, $o_{\text{cloud, 1:t}}$) to generate the training labels described next.
i) Label Generation for RGB-D Encoder $f_{\text{semantic}}$. We generate training labels for the semantic decoder $f_{\text{semantic}}$ by passing $o_{\text{rgb, t}}$ through a frozen Dinov2 encoder and bilinearly interpolating the spatial dimension of our output feature map to match the spatial resolution of $y_{\text{semantic}}$. We generate depth completion labels $\hat{y}_{\text{depth}}$ by projecting $o_{\text{cloud, t}}$ to the image and applying bilateral filtering [28] for edge-aware inpainting, producing a dense depth map that respects object boundaries.
ii) Label Generation for BEV Inpainting Backbone $f_{\text{bev}}$. To train $f_{\text{bev}}$, we first generate $\hat{y}_{\text{bev, dynamic}}$ by prompting SAM2 with bounding boxes of common dynamic classes (e.g., vehicles, pedestrians), backproject the masks to 3D using $o_{\text{cloud, t}}$, and apply DBSCAN [6] clustering at multiple density thresholds. The dynamic BEV map retains clusters with sufficient IoU overlap with the SAM2 instance labels. This procedure mirrors prior works [26] that leverage SAM2 for learning instance-aware 3D representations. To generate $\hat{y}_{\text{bev, static}}$, we query SAM2 with a grid of points, filter out overlaps with dynamic masks to isolate static segments (generating dynamic masks as before), and apply a greedy IoU-based merging strategy across frames to maintain instance consistency across frames. The merged static masks are then projected and accumulated in BEV using known robot poses. We refer readers to the Appendix Sec. X-A for our algorithmic formulation describing the merging procedure. Lastly, we generate $\hat{y}_{\text{bev, elev}}$ by accumulating static 3D points across sequential frames using robot poses (generating static masks as before), assign each 3D point to a grid cell, and compute each cell’s minimum elevation by averaging the N lowest 3D points that fall into that cell.
@@ IV-B2 Training the Reward Function $r_{\phi}$
After freezing $\Theta$, we train $r_{\phi}$ using our counterfactual-based reward learning objective $\mathcal{L_{\text{IRL}}}$, where counterfactuals hold all other variables constant with the expert except for the path taken to the end goal. $\mathcal{L_{\text{IRL}}}$ optimizes the reward function such that the likelihood of counterfactual trajectories is minimized and the likelihood of expert trajectories is maximized, which we observe improves the sample efficiency, expert alignment, and interpretability of the learned rewards.
To understand $\mathcal{L}_{\text{IRL}}$, we formally refer to counterfactuals as suboptimal trajectories and define a state-action visitation distribution to be the discounted probability of reaching a state $s$ under a policy $\pi$ and taking action $a$: $\rho^{\pi}(s,a)=(1-\gamma)\sum\gamma^{t}p(s_{t}=s,a_{t}=a|\pi)$. We define each state $s$ to be an xy location on the local BEV grid and our actions $a$ along the 8 connected grid. Under this formulation, the Counterfactual IRL objective is simple: For each BEV observation $y_{bev}$, given state-actions sampled from the expert’s visitation distribution $\rho^{E}$ and state-actions sampled from the suboptimal visitation distribution $\rho^{S}$; we obtain a reward function and policy by solving the following optimization problem:
@@
where $\rho_{\pi}$ denotes current policy visitation and $\alpha$ is a tunable hyperparameter that balances the relative importance between suboptimal and expert demonstrations.
Using the relationship defined in Eq. 10, we can rewrite Eq. 6 in terms of the state-action visitation distribution to obtain our Counterfactual IRL training objective:
@@
Assuming non-trivial rewards, which can be enforced using gradient regularization techniques [19], the above objective learns a reward function such that the difference between the expert’s return and agent policy’s return is minimized while ensuring suboptimal counterfactuals have a low return. Next, we describe our active learning framework for learning $r_{\phi}$ from counterfactuals and how we generate these counterfactual annotations. We conclude by deriving the reward learning objective using the Bradley-Terry model of preferences and show connections to Inverse Reinforcement Learning (IRL).
Active Reward Learning from Counterfactuals. Our reward learning framework uses Counterfactual IRL to learn a mapping from BEV features in $y_{\text{bev}}$ to their scalar utilities. Our framework trains $r_{\phi}$ in multiple phases, iteratively prompting expert operators for counterfactual annotations in underperforming scenarios until the learned rewards are sufficient for planning trajectories similar to expert demonstrations. Fig. 3 illustrates this framework, which we describe next:
@@
Phase I: Warmstart We obtain a base reward function by training $r_{\phi}$ using only expert demonstrations. This can be done simply by setting $\alpha$ equal to zero in $\mathcal{L}_{\text{IRL}}$.
Phase II: Synthetic Counterfactual Generation We rollout a policy using the learned rewards from phase I and select training samples needing refinement using the Hausdorff distance between the policy rollouts and expert demonstrations. For samples exceeding a distance threshold, we accumulate RGB observations to BEV and generate candidate trajectories sharing start/end points with the expert. A human annotator using our counterfactual annotation tool in Fig. 8 to select trajectories that violate their preferences (e.g., collisions, undesirable terrain) as counterfactuals, which are used in phase III to retrain the reward function using $\mathcal{L}_{\text{IRL}}$.
@@
Phase III: Counterfactual Reward Alignment We retrain $r_{\phi}$ using $\mathcal{L}_{\text{IRL}}$ using phase II’s counterfactual annotations and the original expert demonstrations. We set $\alpha$ to be nonzero to balance their relative importance. We repeat phases II and III until the learned policy converges with expert behavior.
@@
Counterfactual Generation Details. Counterfactual IRL relies on being able to sample counterfactual trajectories that visit a diverse set of states between the start and goal. To do this, we propose a simple method of perturbing the expert trajectory to obtain these counterfactuals. Given the expert trajectory $\Gamma^{E}_{t}$, we sample a handful of ”control” states at regular intervals along $\Gamma^{E}_{t}$, excluding the start and goal states. We perturb each control state according to a non-zero mean Gaussian distribution centered at each control state $\mathcal{N}(\mu+s,\sigma)$ before planning a kinematically feasible path that reaches the start, goal, and perturbed control states. Practically, we implement this using Hybrid A* [5] - however, any kinematic planner will suffice. We repeat this process twice using positive and negative $\mu$ terms to sample paths ($\sim 10$) on both sides of the expert before passing these paths to the operator for counterfactual labeling. For additional implementation details regarding generating alternate trajectories, we refer readers to Appendix Sec. X-C.
@@
Counterfactual IRL Derivation $\mathcal{L}_{\text{IRL}}$. IRL methods [50, 23] allow for learning reward functions $r_{\phi}$, parameterized by $\phi$, given expert demonstrations. However, they provide no mechanism to incorporate suboptimal trajectories. Suboptimal trajectories are easy to obtain and enable a data flywheel for navigation; the BEV observations obtained from expert runs can simply be relabeled with suboptimal or unsafe trajectories offline without any more environmental interactions. Motivated by this idea, we derive $\mathcal{L_{\text{IRL}}}$, a general and principled way to learn from suboptimal and expert trajectories jointly under a single objective.
@@
This property extends to any suboptimal visitations, and as a consequence of linearity of expectations, to any convex combination of the current policy’s visitation and any other suboptimal visitation distribution. Mathematically, this is:
@@
$\min_{\pi}\max_{\phi}\left(J^{E}(r_{\phi})-(\alpha J^{S}(r_{\phi})+(1-\alpha)J^{\pi}(r_{\phi}))\right).$ (14)
@@ V-1 Global and Local Planning and Controls
Our global planning module plans a global path and identifies the next local subgoal for our local planner to plan local paths. Given a user-specified GPS end goal $G_{N}$, we use OpenStreetMap [24] to obtain a semi-dense sequence of coarse GPS goals $G=[G_{1},...,G_{N}]$ spaced 10 meters apart. To select the next GPS subgoal, we compute the set of distances $D=\{||g-G_{N}||\text{ }|\text{ }g\in G\cup\{G_{t}\}\}$, where $D$ contains the distance of the robot $G_{t}$ from $G_{N}$ and the distance of each GPS subgoal in $G$ from $G_{N}$. From the set of subgoals with a smaller distance to $G_{N}$ than $G_{t}$, we select the farthest subgoal to use as the next subgoal. We project this subgoal on the edge of the local planning horizon, a 6 meter circle around the robot, giving us a carrot for local path planning.
We adopt a DWA [8] style approach to local path planning, where we enumerate a set of constant curvature arcs (31 in our case) from the egocentric robot frame. We compute a scalar cost for each trajectory by sampling points along each each arc and computing the discounted cost at each point using our predicted reward map. We simply invert and normalize our reward map to the range $[0,1]$ to convert it to a costmap. Finally, we compute the distance between the last point on the arc with the local carrot to obtain a goal-reaching cost. We multiply the learned and goal-reaching costs by tunable weights before selecting the trajectory with the lowest cost. Finally, we perform 1D time optimal control [27] to generate low-level actions to follow this trajectory. Empirically, we find that sampling 30 points and using a discount factor of 0.95 is sufficiently dense. We find that tuning the goal-reaching cost to be 1/10th the importance of the learned cost achieves good balance between goal-reaching and adhering to operator preferences.
@@
  • ($Q_{3}$) How much do counterfactual demonstrations improve urban navigation performance?